# V2X-M2C: Efficient Multi-Module Collaborative Perception with Two Connections

### v1

@@ 4.3 Qualitative Evaluation
Visualization of Attention Weights. In Figure 6, we visualize the attention weights in A-Att, where the brighter colors indicate the areas the ego agent focuses on. Some observations to notice are as follows: i) Our models show a clear difference between the regions that are paid attention to or not, compared to the V2X-ViT’s attention weights (Figure 6-(b)). ii) Our models have different attention tendencies. For M2C-P in Figure 6-(d), the ego agent relies on its own confidence regions where its point clouds are dense and not occluded. Also, it relies on the aux agent’s confidence regions to incorporate wide-range information. In M2C-S in Figure 6-(c), the ego agent focuses on the areas where objects densely exist. The tendencies are more explicitly identified in the infrastructure data, which has a wider detection range but is sparse and can be noisy at far distances. The ego agent in M2C-S leverages the infrastructure data heavily, simultaneously differentiating between roads and objects. On the other hand, the ego agent in M2C-P takes the infrastructure data relatively less because the infrastructure data has lower confidence than the data from nearby aux agents.
4.4 Ablation Study.
Effect of sub-modules. We present the ablation study results in Table 4. The inter-agent complementary information, which is related to A-Att, affects both models significantly. In the perfect/simple noise setting, the performances of M2C-S decrease by 20.25%/12.47%, and that of M2C-P decreases by 18.09%/10.22% at AP@0.7 by eliminating A-Att. H-Conv also affects M2C-S and M2C-P, which decreases the performances by 2.78%/4.96% and 1.87%/4.22%, respectively.
Table 2: Effect of sub-modules in V2X-M2C
A-Att S-Att H-Conv Original Perfect Simple Noise
AP@0.5/0.7 AP@0.5/0.7
M2C-S - 0.805 / 0.684 0.809 / 0.683
- 0.910 / 0.851 0.899 / 0.768
- 0.883 / 0.835 0.864 / 0.743
- 0.922 / 0.860 0.910 / 0.781
M2C-P 0.806 / 0.698 0.808 / 0.698
0.893 / 0.849 0.883 / 0.764
0.883 / 0.836 0.863 / 0.745
0.898 / 0.847 0.893 / 0.776
0.904 / 0.853 0.894 / 0.785
Table 3: Effect of comp. ratio at AP@0.7 in M2C-S. Noisy means the simple noise.
Rate Perfect Noisy # Params GFLOPs
$\times$1 0.864 0.800 24.76M 790
$\times$2 0.865 0.798 9.75M 230
$\times$4 (Ours) 0.860 0.781 5.94M 103
$\times$8 0.842 0.751 4.97M 74
Table 4: Ablation study of HRPE at AP@0.7. Noisy means the simple noise.
Model HRPE Perfect Noisy
M2C-S 0.860 0.781
0.858 0.765
M2C-P 0.853 0.785
0.855 0.772
CCL compression ratio. We train M2C-S with various compression rates in CCL to validate whether the reduction in feature size causes information loss. Table 4 shows that the rate of $\times$2 yields the best performance at AP@0.7. Although the performance of the rate of $\times$4 is lower by 2.24% compared with the rate of $\times$2 in the perfect and noisy settings, the number of parameters and GFLOPs are decreased by 39.1% and 55.22%, respectively. From the efficiency perspective, we select the rate of $\times$4 as the final setting.
Effect of HRPE. We use HRPE in A-Att to help the module understand the features of the other agents. Table 4 shows that HRPE significantly enhances the performance of the two models in the noise setting, while it does not contribute much in the perfect setting. The results show that HRPE, by setting the two hyperparameters $\theta$ and $d$ (given in Appendix B.3), makes our models robust to noise.
5 Conclusion
This paper has proposed efficient collaborative perception models, M2C-S and M2C-P, based on two different connection types. The two critical components that enhance the effectiveness and efficiency of our models are CCL and H-Conv. CCL reduces the feature size and makes our fusion module lightweight, and H-Conv handles multiple heterogeneous agents and simultaneously prevents our models from being computationally heavy. In addition, we have compared and analyzed M2C-S and M2C-P through extensive experiments regarding detection accuracy, noise robustness, and learning tendency. We envision these analyses providing important guidelines to design new model architectures for collaborative perception considering different connections of multiple features.
Limitation and future work. Although we have analyzed the characteristics of our two models and achieved the SOTA performances, we did not propose a single integrated model that shows superior performance in all aspects. In future work, we plan to integrate the parallel and sequential models into a single architecture and improve A-Att and S-Att along with the integration.
Appendix A Our Contributions
  • We propose the collaborative perception model, V2X-M2C for 3D object detection. V2X-M2C is divided into two models: M2C-S connects the sub-modules in series, and M2C-P connects them in parallel. We consider multiple heterogeneous agents, making our model expandable to V2X scenarios. Our two models achieve state-of-the-art (SOTA) effectiveness, efficiency, and robustness to noise.
  • We perform the qualitative comparison between the two connection types in terms of how each sub-module influences the each other.
  • We utilize a channel compression layer (CCL) to make the overall fusion module lightweight, thereby significantly enhancing model efficiency while maintaining performance.
  • We design a heterogeneity-reflected convolution module (H-Conv) to consider multiple heterogeneous agents with a low number of parameters and low computation complexity.
Appendix B Implementation Details
B.1 Dataset.
We use the V2XSet [2] and OPV2V [18] datasets to train and validate models in both V2V and V2X scenarios.
V2XSet. V2XSet is a simulated dataset supporting V2X perception, co-simulated using CARLA [19] and OpenCDA [20]. It comprises 73 scenes with a minimum of 2 to 5 connected agents and includes 11K 3D annotated LiDAR point cloud frames. The training, validation, and testing sets consist of 6.7K, 2K, and 2.8K frames, respectively.
OPV2V. OPV2V is a simulated dataset designed for multi-agent V2V perception. Each frame typically comprises approximately 3 CAVs, with a minimum of 2 and a maximum of 7. It includes 10.9K LiDAR point cloud frames with 3D annotations. The training, validation, and testing splits include 6.8K, 2K, and 2.2K frames, respectively.
B.2 Comparison Models.
We select comparison models through several considerations to check whether our models are significant and reliable: i) The model has been suggested within the past two years. ii) The model has been cited in many related works. iii) The model has distinct methods or characteristics from each other. Considering these factors, we select CoBEVT [1], V2X-ViT [2], and Where2comm [3]. Unlike other models, CoBEVT presents a framework for multi-agent multi-camera perception that collaboratively produces BEV map predictions using both camera and LiDAR. V2X-ViT attempts to utilize heterogeneous agents, such as vehicles and road infrastructure, and solve the problem of heterogeneity between them using a transformer with a graph structure. Where2comm introduces a spatial confidence map, enabling agents to share only spatially sparse yet perceptually important information.
B.3 Hyperparameter Details.
Training Parameters. For a fair comparison, we train both our models and SOTA models using AdamW [21], with a learning rate of 3e-4 and a weight decay of 0.01. We use the same learning scheduler, Cosine Annealing Warm-Up Restarts [34], applying a warm-up learning rate of 2e-4 and warm-up epochs. We train the models for up to 40 epochs and use a maximum of five connected agents, including the ego agent. The models are trained and validated on an RTX 4090.
Model Parameters. Regarding the fusion module, HRPE in A-att utilizes the relative angle $\theta$ and distance $d$. In M2C-S, the relative angle and distance are segmented into 10°10°10  ° and 30 m, respectively. In M2C-P, these are divided into 10°10°10  ° and 20 m. HRPE then uses the integer value rounded from the divided terms. In S-att, the lightweight DiNAT [17] features two depths. For DiNAT’s hyperparameters, we use a $7\times 7$ kernel with dilation rates of 4 and 2 applied to each depth, respectively. The H-conv comprises three blocks, each incorporating heterogeneity-reflected convolution using a $3\times 3$ kernel. The first and second blocks employ a residual connection linking each block’s input and output, while the third block does not.
B.4 Training Details.
We train the models in three different noise settings: 1) perfect setting, 2) simple noise setting, and 3) harsh noise setting. Each noise setting differs in the compression ratio for feature sharing and noise values $(t_{\mathrm{lag}},\sigma_{\mathrm{loc}},\sigma_{\mathrm{hdg}})$.
Perfect Setting. Models are trained up to 40 epochs, and the detailed configuration is as follows.
  • Compression Ratio: ×$\times 32$
  • Noise Value: (0 ms, 0.0°0.0°0.0  °, 0.0 m)
Simple Noise Setting. Models are trained up to 40 epochs, and the detailed configuration is as follows.
  • Compression Ratio: ×$\times 32$
  • Noise Value: (constant 100 ms, 0.2°0.2°0.2  °, 0.2 m)
Harsh Noise Setting. Training in the harsh noise setting consists of three steps: 1) perfect setting training, 2) compressed setting fine-tuning, and 3) noise setting fine-tuning.
The perfect setting training trains the models up to 40 epochs with the configurations below.
  • Compression Ratio: ×$\times 1$ (not compressed)
  • Noise Value: (0 ms, 0.0°0.0°0.0  °, 0.0 m)
After completing the perfect setting training, we select the model with the best performance weight from the 31st to 40th epoch at AP@0.7. Then, we train the model for an additional two epochs. At this step, we convert the compression ratio from ×$\times 1$ to ×$\times 32$. Finally, we train the model for an additional thirty-eight epochs with the noise settings as below.
  • Compression Ratio: ×$\times 32$
  • Noise Value: (random 200 ms, 0.2°0.2°0.2  °, 0.2 m)
B.5 Inference Details.
The results of training differ even when the model has the same architecture and hyperparameters due to the properties of gradient descent. Additionally, the random seed influences the probability distribution, and this distribution affects the difficulty level of scenarios. To ensure model reliability, we evaluate models through several inferences depending on the noise setting: 1) perfect setting, 2) simple noise setting, and 3) harsh noise setting.
Perfect Setting. In the perfect setting, we utilize three different pre-trained models that have been trained under this setting. We evaluate each model’s performance by averaging the AP@0.5/0.7 of thirty weights that are derived from the pre-trained models during the 31st to 40th epochs.
Simple Noise Setting. Regarding the simple noise setting, we consider noise variations by seed. We used nine models pre-trained under the simple noise setting, applying three different seeds (25, 26, and 27). The process is the same as in the perfect setting, except for evaluating the performance of the ablation study. For the ablation study’s performance evaluation, we simplify the process by only applying seed 27. We use seed 27 because its performance demonstrates typical results upon observation; seed 25 results in easy noise, and seed 26 in hard noise.
Harsh Noise Setting. To evaluate noise robustness, we use models pre-trained with harsh noise settings under three seeds (25, 26, and 27). We calculate AP@0.5/0.7 depending on three harsh noise components: random latency, localization, and heading noise standard deviation. We create sixteen noise sets by varying one component while setting the others to zero: random $t_{\mathrm{lag}}\in[0\,\mathrm{ms},500\,\mathrm{ms}]$, $\sigma_{\mathrm{loc}}\in[0.0\,\mathrm{m},0.5\,\mathrm{m}]$, and $\sigma_{\mathrm{hdg}}\in[$$,$$]$ 0.0  ° , 1.0  ° ], including the zero noise scenario (0 ms, 0.0 m, 0.0°0.0°0.0  °). We use the weight from the pre-trained model results from the 36th to 40th epochs and calculate AP@0.5/0.7 for each weight depending on each noise set. Finally, we evaluate the performance by averaging AP@0.5/0.7 for each noise set.
Table 5: Robustness to various ranges of noises with detailed values.
Model Random Latency ($t_{\mathrm{lag.}}$)
@@ Appendix B Implementation Details
CoBEVT [1] 0.767 0.761 0.742 0.709 0.687 0.672
V2X-ViT [2] 0.772 0.765 0.752 0.727 0.696 0.685
Where2comm [3] 0.776 0.768 0.750 0.725 0.708 0.699
M2C-S (Ours) 0.776 0.770 0.760 0.736 0.716 0.707
M2C-P (Ours) 0.788 0.781 0.771 0.745 0.725 0.717
Model Location Noise std. ($\sigma_{\mathrm{loc.}}$)
@@
CoBEVT [1] 0.767 0.765 0.743 0.712 0.678 0.656
V2X-ViT [2] 0.772 0.775 0.763 0.738 0.703 0.671
Where2comm [3] 0.776 0.775 0.755 0.725 0.695 0.670
M2C-S (Ours) 0.776 0.781 0.772 0.746 0.710 0.679
M2C-P (Ours) 0.788 0.793 0.783 0.753 0.718 0.687
Model Heading Error std. ($\sigma_{\mathrm{hdg.}}$)
@@
CoBEVT [1] 0.767 0.769 0.758 0.738 0.717 0.701
V2X-ViT [2] 0.772 0.775 0.771 0.759 0.747 0.737
Where2comm [3] 0.776 0.776 0.769 0.757 0.745 0.733
M2C-S (Ours) 0.776 0.781 0.779 0.767 0.754 0.742
M2C-P (Ours) 0.788 0.794 0.792 0.781 0.769 0.755
Table 6: Robustness to a subset of noises with detailed values.
Model Random Latency ($t_{\mathrm{lag.}}$)
@@
CoBEVT [1] 0.767 0.751 (1.99% $\downarrow$) 0.690 (10.06% $\downarrow$)
V2X-ViT [2] 0.772 0.758 (1.80% $\downarrow$) 0.703 (9.01% $\downarrow$)
Where2comm [3] 0.776 0.759 (2.21% $\downarrow$) 0.711 (8.42% $\downarrow$)
M2C-S (Ours) 0.776 0.765 (1.39% $\downarrow$) 0.720 (7.21% $\downarrow$)
M2C-P (Ours) 0.788 0.776 (1.55% $\downarrow$) 0.729 (7.54% $\downarrow$)
Model Location Noise std. ($\sigma_{\mathrm{loc.}}$)
0.0 m [0.1 ms, 0.2 ms] [0.3 ms, 0.5 ms]
CoBEVT [1] 0.767 0.754 (1.61% $\downarrow$) 0.682 (11.04% $\downarrow$)
V2X-ViT [2] 0.772 0.769 (0.44% $\downarrow$) 0.704 (8.81% $\downarrow$)
Where2comm [3] 0.776 0.765 (1.42% $\downarrow$) 0.697 (10.27% $\downarrow$)
M2C-S (Ours) 0.776 0.777 (0.13% $\uparrow$) 0.711 (8.30% $\downarrow$)
M2C-P (Ours) 0.788 0.788 (0.03% $\uparrow$) 0.719 (8.75% $\downarrow$)
@@
CoBEVT [1] 0.767 0.763 (0.43% $\downarrow$) 0.719 (6.23% $\downarrow$)
V2X-ViT [2] 0.772 0.773 (0.12% $\uparrow$) 0.748 (3.17% $\downarrow$)
Where2comm [3] 0.776 0.773 (0.44% $\downarrow$) 0.745 (4.02% $\downarrow$)
M2C-S (Ours) 0.776 0.780 (0.60% $\uparrow$) 0.754 (2.79% $\downarrow$)
M2C-P (Ours) 0.788 0.793 (0.59% $\uparrow$) 0.768 (2.52% $\downarrow$)
Appendix C Experiments.
C.1 Noise Robustness.
To evaluate the robustness against noise, we train and validate models under the harsh noise setting (Sections B.4 and B.5). As shown in Table 5 with detailed values, our M2C-S and M2C-P outperform the SOTA models across all noise components. To present a detailed comparison, we divide the noise sets into two subsets: 1) normal noise and 2) severe noise. The normal noise subset is based on the normal range defined in the work [2]. In this paper, we define a normal noise subset as $t_{\mathrm{lag.}}\in[100\,\mathrm{ms},200\,\mathrm{ms}]$, $\sigma_{\mathrm{loc.}}\in[0.1\,\mathrm{m},0.2\,\mathrm{m}]$, and $\sigma_{\mathrm{hdg.}}\in[$$,$$]$ 0.2  ° , 0.4  ° ], and the severe noise subset as $t_{\mathrm{lag.}}\in[300\,\mathrm{ms},500\,\mathrm{ms}]$, $\sigma_{\mathrm{loc.}}\in[0.3\,\mathrm{m},0.5\,\mathrm{m}]$, and $\sigma_{\mathrm{hdg.}}\in[$$,$$]$ 0.6  ° , 1.0  ° ]. In Table 6, we evaluate performance by averaging each component in the subset.
Normal Noise. Under normal noise conditions, our models outperform the SOTA models. In particular, M2C-P shows the best performance, improving performance in terms of latency, localization noise, and heading noise by 2.23%, 2.56%, and 2.57% compared to the highest performance. Moreover, both M2C-S and M2C-P demonstrate the lowest sensitivity to noise. Our models exhibit performance enhancements regarding localization and heading noise, while most SOTA models show reduced performance.
Severe Noise. Our models outperform the SOTA models in severe noise conditions. M2C-P increases by 2.23%/2.56%/2.57% compared to the highest performance. In addition, our models show the lowest sensitivity to severe noise.
C.2 Detection Visualization.
We visualize the detection results in the V2V and V2I scenarios, respectively. In the V2V scenario (Fig. 7), our models demonstrate robustness to occlusion. Moreover, the bounding boxes generated by our models align well with the detected objects, showing superior alignment compared to those produced by the SOTA models. In the V2I scenario (Fig. 8), our models effectively utilize sparse point clouds from the infrastructure for detection. Additionally, the predicted bounding boxes are more precise than those of the SOTA models.
C.3 Attention Weights Visualization.
In Figure 9, we visualize additional the attention weights in agent-wise attention of models in V2V scenario, where the brighter colors indicate the areas the ego agent focuses on. Some observations to notice are as follows: i) Our models show a clear difference between the regions that are paid attention to or not, compared to the V2X-ViT’s attention weights (Figure 9-(b)). ii) Our models have different attention tendencies. For the M2C-P in Figure 9-(d), the ego agent relies on its own confidence regions where its point clouds are dense and not occluded and also on the aux agent’s confidence regions to incorporate wide-range information. In the M2C-S in Figure 9-(c), the ego agent focuses on the areas where objects densely exist.
Figure 7: Detection visualization of V2V scenarios. The green bounding boxes represent the ground truth, indicating vehicles driving on the road, while the red bounding boxes denote predictions. These V2V scenarios are extracted from the V2XSet and OPV2V datasets.
Figure 8: Detection visualization of V2I scenarios. The green bounding boxes represent the ground truth, indicating vehicles driving on the road, while the red bounding boxes denote predictions. These V2I scenarios are extracted from the V2XSet datasets.
(a) V2V Scenario      (b) V2X-ViT (V2V)  (c) Sequential (V2V)   (d) Parallel (V2V)
Figure 9: Aggregated LiDAR points and attention map for V2V. The white circle is the ego agent, and the blue and green circles are aux agents. Small grey circles are vehicles in the detection range.

### v2

@@ Appendix A Appendix
A.5 Noise Robustness.
As shown in Table 12, 13, 14, and 18, we show the detailed values used in Section 4.2, which is ParCon outperforms the SOTA models across all noise ranges.
To support our claims of noise robustness, we analyze an additional comparison, and then divide the noise set into two subsets: 1) normal noise and 2) severe noise. The normal noise subset is based on the normal range defined in the work Xu et al. [2022b]. Hence, we define a normal noise subset as $t_{\mathrm{lag.}}\in[100\,\mathrm{ms},200\,\mathrm{ms}]$, $\sigma_{\mathrm{loc.}}\in[0.1\,\mathrm{m},0.2\,\mathrm{m}]$, and $\sigma_{\mathrm{hdg.}}\in[$$,$$]$ 0.2  ° , 0.4  ° ], and the severe noise subset as $t_{\mathrm{lag.}}\in[300\,\mathrm{ms},500\,\mathrm{ms}]$, $\sigma_{\mathrm{loc.}}\in[0.3\,\mathrm{m},0.5\,\mathrm{m}]$, and $\sigma_{\mathrm{hdg.}}\in[$$,$$]$ 0.6  ° , 1.0  ° ]. In Table 15, 16, 17, and 19, we evaluate performance by averaging each component in the subset.
Normal Noise. Under normal noise conditions, ParCon outperforms SOTA models in terms of latency, heading, and localization noise. Moreover, ParCon demonstrates the lowest sensitivity to noise, decreasing by 0.03%, 0.18%, and 1.81& on V2XSet, 1.23%, 0.77%, and 2.28% on OPV2V under latency/heading noise/localization noise, and 0.92%, 3.49% on DAIR-V2X under heading/localization noise, repectively. Our models exhibit performance enhancements regarding localization and heading noise, while most SOTA models show reduced performance.
Severe Noise. Our models always outperform the SOTA models in severe noise conditions. Also, our models show the lowest sensitivity to severe noise. ParCon has low drop rate by 4.01%/2.13%/7.75% on V2XSet and by 3.38%/3.32%/10.29% on OPV2V under latency/heading noise/localization noise, respectively. Although the drop rate of our method is not always the lowest in some dataset, ParCon always keeps SOTA performance.
Table 12: Robustness to various ranges of noises with detailed values on V2XSet.
Model Maximum Latency ($t_{\mathrm{lag.}}$)
@@
CoBEVT 0.764 0.759 0.738 0.712 0.696 0.684
V2X-ViT 0.747 0.738 0.721 0.701 0.683 0.673
Where2comm 0.741 0.741 0.722 0.697 0.688 0.679
CoAlign 0.718 0.711 0.695 0.677 0.670 0.665
ParCon (Ours) 0.760 0.765 0.754 0.737 0.729 0.723
Model Heading Noise std. ($\sigma_{\mathrm{loc.}}$)
CoBEVT 0.764 0.766 0.760 0.750 0.735 0.721
V2X-ViT 0.747 0.748 0.745 0.739 0.731 0.721
Where2comm 0.741 0.740 0.731 0.720 0.710 0.701
CoAlign 0.718 0.717 0.712 0.705 0.698 0.691
ParCon (Ours) 0.760 0.760 0.757 0.751 0.744 0.737
Model Localization Error std. ($\sigma_{\mathrm{hdg.}}$)
@@
CoBEVT 0.764 0.752 0.711 0.634 0.545 0.465
V2X-ViT 0.747 0.737 0.717 0.693 0.668 0.645
Where2comm 0.741 0.730 0.700 0.661 0.628 0.605
CoAlign 0.718 0.703 0.671 0.632 0.599 0.573
ParCon (Ours) 0.760 0.754 0.738 0.719 0.700 0.683
Table 13: Robustness to various ranges of noises with detailed values on OPV2V.
Model Maximum Latency ($t_{\mathrm{lag.}}$)
0 ms 100 ms 200 ms 300 ms 400 ms 500 ms
CoBEVT 0.659 0.656 0.646 0.630 0.612 0.593
V2X-ViT 0.685 0.682 0.677 0.670 0.664 0.658
Where2comm 0.675 0.673 0.661 0.647 0.632 0.621
CoAlign 0.638 0.634 0.624 0.614 0.603 0.593
ParCon (Ours) 0.713 0.711 0.705 0.697 0.690 0.683
@@
CoBEVT 0.659 0.656 0.646 0.630 0.612 0.593
V2X-ViT 0.685 0.682 0.677 0.670 0.664 0.658
Where2comm 0.675 0.673 0.661 0.647 0.632 0.621
CoAlign 0.638 0.634 0.624 0.614 0.603 0.593
ParCon (Ours) 0.713 0.711 0.705 0.697 0.690 0.683
Model Localization Error std. ($\sigma_{\mathrm{hdg.}}$)
0.0 m 0.1 m 0.2 m 0.3 m 0.4 m 0.5 m
CoBEVT 0.659 0.639 0.575 0.490 0.416 0.351
V2X-ViT 0.685 0.679 0.575 0.490 0.416 0.351
Where2comm 0.675 0.661 0.615 0.569 0.536 0.510
CoAlign 0.638 0.621 0.569 0.517 0.480 0.456
ParCon (Ours) 0.713 0.706 0.688 0.663 0.639 0.618
Table 14: Robustness to various ranges of noises with detailed values on DAIR-V2X.
Model Heading Noise std. ($\sigma_{\mathrm{loc.}}$)
CoBEVT 0.448 0.446 0.440 0.433 0.427 0.421
V2X-ViT 0.449 0.447 0.441 0.434 0.428 0.423
Where2comm 0.404 0.404 0.400 0.396 0.391 0.387
CoAlign 0.408 0.407 0.403 0.399 0.394 0.391
ParCon (Ours) 0.466 0.464 0.458 0.452 0.445 0.440
Model Localization Error std. ($\sigma_{\mathrm{hdg.}}$)
0.0 m 0.1 m 0.2 m 0.3 m 0.4 m 0.5 m
CoBEVT 0.448 0.439 0.424 0.408 0.397 0.388
V2X-ViT 0.449 0.439 0.426 0.411 0.401 0.394
Where2comm 0.404 0.398 0.387 0.376 0.368 0.363
CoAlign 0.408 0.402 0.393 0.384 0.376 0.371
ParCon (Ours) 0.466 0.458 0.441 0.426 0.414 0.406
Table 15: Robustness to a subset of noises with detailed values on V2XSet.
Model Maximum Latency ($t_{\mathrm{lag.}}$)
@@
CoBEVT 0.764 0.748 (2.09% $\downarrow$) 0.697 (8.77% $\downarrow$)
V2X-ViT 0.747 0.730 (2.24% $\downarrow$) 0.686 (8.16% $\downarrow$)
Where2comm 0.741 0.731 (1.35% $\downarrow$) 0.688 (7.21% $\downarrow$)
CoAlign 0.718 0.703 (2.09% $\downarrow$) 0.671 (6.54% $\downarrow$)
ParCon (Ours) 0.760 0.760 (0.03% $\downarrow$) 0.729 (4.01% $\downarrow$)
@@
CoBEVT 0.764 0.763 (0.17% $\downarrow$) 0.735 (3.79% $\downarrow$)
V2X-ViT 0.747 0.746 (0.05% $\downarrow$) 0.730 (2.17% $\downarrow$)
Where2comm 0.741 0.735 (0.79% $\downarrow$) 0.710 (4.16% $\downarrow$)
CoAlign 0.718 0.714 (0.47% $\downarrow$) 0.698 (2.75% $\downarrow$)
ParCon (Ours) 0.760 0.759 (0.18% $\downarrow$) 0.744 (2.13% $\downarrow$)
Model Localization Noise std. ($\sigma_{\mathrm{loc.}}$)
0.0 m [0.1 ms, 0.2 ms] [0.3 ms, 0.5 ms]
CoBEVT 0.764 0.732 (4.26% $\downarrow$) 0.548 (28.28% $\downarrow$)
V2X-ViT 0.747 0.727 (2.58% $\downarrow$) 0.669 (10.43% $\downarrow$)
Where2comm 0.741 0.715 (3.61% $\downarrow$) 0.631 (14.82% $\downarrow$)
CoAlign 0.718 0.687 (4.26% $\downarrow$) 0.601 (16.20% $\downarrow$)
ParCon (Ours) 0.760 0.746 (1.81% $\downarrow$) 0.701 (7.75% $\downarrow$)
Table 16: Robustness to a subset of noises with detailed values on OPV2V.
Model Maximum Latency ($t_{\mathrm{lag.}}$)
0 ms [100 ms, 200 ms] [300 ms, 500 ms]
CoBEVT 0.659 0.644 (2.33% $\downarrow$) 0.610 (7.43% $\downarrow$)
V2X-ViT 0.685 0.662 (3.37% $\downarrow$) 0.636 (7.13% $\downarrow$)
Where2comm 0.675 0.661 (2.15% $\downarrow$) 0.642 (4.97% $\downarrow$)
CoAlign 0.638 0.626 (1.99% $\downarrow$) 0.611 (4.26% $\downarrow$)
ParCon (Ours) 0.713 0.705 (1.23% $\downarrow$) 0.689 (3.38% $\downarrow$)
Model Heading Noise std. ($\sigma_{\mathrm{hdg.}}$)
CoBEVT 0.659 0.651 (1.24% $\downarrow$) 0.612 (7.17% $\downarrow$)
V2X-ViT 0.685 0.679 (0.85% $\downarrow$) 0.664 (3.06% $\downarrow$)
Where2comm 0.675 0.667 (1.27% $\downarrow$) 0.633 (6.28% $\downarrow$)
CoAlign 0.638 0.629 (1.45% $\downarrow$) 0.603 (5.52% $\downarrow$)
ParCon (Ours) 0.713 0.708 (0.77% $\downarrow$) 0.690 (3.32% $\downarrow$)
Model Localization Noise std. ($\sigma_{\mathrm{loc.}}$)
0.0 m [0.1 ms, 0.2 ms] [0.3 ms, 0.5 ms]
CoBEVT 0.659 0.607 (7.89% $\downarrow$) 0.419 (36.40% $\downarrow$)
V2X-ViT 0.685 0.670 (2.24% $\downarrow$) 0.608 (11.30% $\downarrow$)
Where2comm 0.675 0.638 (5.53% $\downarrow$) 0.538 (20.32% $\downarrow$)
CoAlign 0.638 0.595 (6.74% $\downarrow$) 0.484 (24.12% $\downarrow$)
ParCon (Ours) 0.713 0.697 (2.28% $\downarrow$) 0.640 (10.29% $\downarrow$)
Table 17: Robustness to a subset of noises with detailed values on DAIR-V2X.
Model Heading Noise std. ($\sigma_{\mathrm{hdg.}}$)
CoBEVT 0.448 0.443 (1.01% $\downarrow$) 0.427 (4.56% $\downarrow$)
V2X-ViT 0.449 0.444 (1.04% $\downarrow$) 0.429 (4.49% $\downarrow$)
Where2comm 0.404 0.402 (0.50% $\downarrow$) 0.391 (3.12% $\downarrow$)
CoAlign 0.408 0.405 (0.78% $\downarrow$) 0.394 (3.37% $\downarrow$)
ParCon (Ours) 0.466 0.461 (0.92% $\downarrow$) 0.446 (4.27% $\downarrow$)
Model Localization Noise std. ($\sigma_{\mathrm{loc.}}$)
0.0 m [0.1 ms, 0.2 ms] [0.3 ms, 0.5 ms]
CoBEVT 0.448 0.431 (3.68% $\downarrow$) 0.398 (11.13% $\downarrow$)
V2X-ViT 0.449 0.432 (3.62% $\downarrow$) 0.402 (10.37% $\downarrow$)
Where2comm 0.404 0.392 (2.89% $\downarrow$) 0.369 (8.63% $\downarrow$)
CoAlign 0.408 0.398 (2.59% $\downarrow$) 0.377 (7.34% $\downarrow$)
ParCon (Ours) 0.466 0.449 (3.49% $\downarrow$) 0.415 (10.75% $\downarrow$)
Table 18: Comparison robustness to various ranges of noises between ParCon-S and ParCon on V2XSet.
Model Maximum Latency ($t_{\mathrm{lag.}}$)
0 ms 100 ms 200 ms 300 ms 400 ms 500 ms
ParCon-S 0.752 0.757 0.746 0.728 0.718 0.711
ParCon (Ours) 0.760 0.765 0.754 0.737 0.729 0.723
Model Heading Noise std. ($\sigma_{\mathrm{loc.}}$)
ParCon-S 0.752 0.752 0.748 0.741 0.734 0.726
ParCon (Ours) 0.760 0.760 0.757 0.751 0.744 0.737
Model Localization Error std. ($\sigma_{\mathrm{hdg.}}$)
0.0 m 0.1 m 0.2 m 0.3 m 0.4 m 0.5 m
ParCon-S 0.752 0.744 0.729 0.709 0.700 0.683
ParCon (Ours) 0.760 0.754 0.738 0.719 0.700 0.683
Table 19: Comparison robustness to a subset of noises between ParCon-S and ParCon with detailed values on V2XSet.
Model Maximum Latency ($t_{\mathrm{lag.}}$)
0 ms [100 ms, 200 ms] [300 ms, 500 ms]
ParCon-S 0.752 0.752 (0.05% $\downarrow$) 0.719 (4.38% $\downarrow$)
ParCon (Ours) 0.760 0.760 (0.03% $\downarrow$) 0.729 (4.01% $\downarrow$)
Model Heading Noise std. ($\sigma_{\mathrm{hdg.}}$)
ParCon-S 0.752 0.750 (0.28% $\downarrow$) 0.734 (2.45% $\downarrow$)
ParCon (Ours) 0.760 0.759 (0.18% $\downarrow$) 0.744 (2.13% $\downarrow$)
Model Localization Noise std. ($\sigma_{\mathrm{loc.}}$)
0.0 m [0.1 ms, 0.2 ms] [0.3 ms, 0.5 ms]
ParCon-S 0.752 0.737 (2.05% $\downarrow$) 0.689 (8.43% $\downarrow$)
ParCon (Ours) 0.760 0.746 (1.81% $\downarrow$) 0.701 (7.75% $\downarrow$)